# Magic Boundaries of 3D Color Codes

### v3

@@ IV.5.3 The $\{y_{\mathbf{z}},g_{\mathbf{z}},p_{\mathbf{z}}\}$-boundary
IV.5.4 The folded boundaries
In this subsection, we investigate the condensation properties of the folded boundaries introduced in Sec. III.3.3.
[…]
@@ IV.5.4 The folded boundaries
[…]
@@
[…]
IV.5.5 The $pg_{\mathbf{x}}py_{\mathbf{x}}yg_{\mathbf{x}}$- and the $y_{\mathbf{z}}g_{\mathbf{z}}p_{\mathbf{z}}$-boundaries
Consider the lattice model we introduce in Sec. III.3.4 and III.3.5. After the conjugation of transversal-$T$ gate, the boundary condensation set becomes $\{y_{\mathbf{z}}g_{\mathbf{z}},g_{\mathbf{z}}p_{\mathbf{z}},pg_{\mathbf{xs}}py_{\mathbf{xs}}yg_{\mathbf{xs}}\}$ and $\{pg_{\mathbf{xs}}py_{\mathbf{xs}},py_{\mathbf{xs}}yg_{\mathbf{xs}},y_{\mathbf{z}}g_{\mathbf{z}}p_{\mathbf{z}}\}$, respectively.
@@
$\{y_{\mathbf{z}}g_{\mathbf{z}},g_{\mathbf{z}}p_{\mathbf{z}},pg_{\mathbf{xs}}py_{\mathbf{xs}}yg_{\mathbf{xs}}\}$ (98)
@@
IV.6 The nested boundaries
Figure 9: An illustration of the $S$ $(\mathrm{CZ}$)-domain wall sweeping across a codimension-1 submanifold in the presence of an boundary with at least one $Z$-type condensation. (a) When the boundary has one $Z$-type condensations, for example $y_{\mathbf{z}}$, the $S$-domain wall on the $pg$-membrane cannot condense on it. (b) However, the other two $S$-domain walls on the $py$- and $yg$-membranes can condense on this boundary. (c) When the boundary has two or more $Z$-type condensations, all of the three types of $S$-domain walls can condense on it.
[…]
IV.6.1 The $\mathbf{c}_{i}$-color boundary
Consider the $\{(\mathbf{c}_{i})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}},(\mathbf{c}_{i}\mathbf{c}_{k})_{\mathbf{x}}\}$-boundary, the $S_{\mathbf{c}_{i}\mathbf{c}_{j}}$- or $S_{\mathbf{c}_{i}\mathbf{c}_{k}}$-domain wall can condense on the boundary. However, the $S_{\mathbf{c}_{j}\mathbf{c}_{k}}$-domain wall can not condense on it. It attaches to the boundary and generates a codimension-2 (1D) domain-wall defect. The condensation set on that defect becomes $\{(\mathbf{c}_{i})_{\mathbf{z}},(\mathbf{c}_{k})_{\mathbf{z}}(\mathbf{c}_{i}\mathbf{c}_{k})_{\mathbf{x}},(\mathbf{c}_{j})_{\mathbf{z}}(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}\}$, while other part of the boundary is unchanged. The corresponding condensation in the TQFT description is $(e_{i},m_{j}e_{k},m_{k}e_{j})$, which has been introduced in Sec. IV.2. The total number of distinct types of the $\mathbf{c}_{i}$-color boundaries with non-condensing $S$-domain wall is 3. We denote this type of boundary as the nested boundary, as illustrated in Fig. 8, Fig. 9(a), and Fig. 10(d).
IV.6.2 The $X$-boundary
@@
[…]
@@
IV.6.4 The $y_{\mathbf{z}}g_{\mathbf{z}}p_{\mathbf{z}}$-boundary
The condensation set on the $y_{\mathbf{z}}g_{\mathbf{z}}p_{\mathbf{z}}$-boundary is $\{pg_{\mathbf{x}}py_{\mathbf{x}},py_{\mathbf{x}}yg_{\mathbf{x}},y_{\mathbf{z}}g_{\mathbf{z}}p_{\mathbf{z}}\}$. The lattice construction is given in Sec. III.3.5. One can show that by sweeping the $S$-domain wall across the codimension-1 submanifold, at the intersection with the boundary, the $S$-domain wall cannot terminate. In the TQFT description, this relation corresponds to Eq. (59). The total number of distinct types of nested boundaries associated with this boundary is 1, as illustrated in Fig. 10(b).
IV.6.5 The other boundaries
@@
IV.7 Summary of boundary types
In summary, we identify 101 distinct types of boundaries for the 3D color code, with 1 of them being the magic boundary and 70 of them being the nested boundaries. While a trivial color permutation ($y\to g$, $g\to p$, $p\to r$, and $r\to y$) can generate additional boundary types, the new boundaries created by this permutation are physically equivalent. Therefore, we exclude these from our count. We summarize the properties of these boundaries in Table 1, Table 2, and Fig. 10.
@@ IV.7 Summary of boundary types
$\{y_{\mathbf{z}},g_{\mathbf{z}},p_{\mathbf{z}}\}$ $(e_{1},e_{2},e_{3})$ 1 Elementary III.1
$\{pg_{\mathbf{x}},py_{\mathbf{x}},yg_{\mathbf{x}}\}$ $(m_{1},m_{2},m_{3})$ 1 Elementary III.2
$\{(\mathbf{c}_{j}\mathbf{c}_{k})_{\mathbf{x}}(\mathbf{c}_{j})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{k})_{\mathbf{x}}(\mathbf{c}_{i})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}\}$ $(m_{i}e_{j},m_{j}e_{i},m_{k})$ 3 Elementary III.3.6
$\{(\mathbf{c}_{j}\mathbf{c}_{k})_{\mathbf{x}}(\mathbf{c}_{j})_{\mathbf{z}}(\mathbf{c}_{k})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{k})_{\mathbf{x}}(\mathbf{c}_{i})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}(\mathbf{c}_{i})_{\mathbf{z}}\}$ $(m_{i}e_{j}e_{k},m_{j}e_{i},m_{k}e_{i})$ 3 Elementary III.3.6
$\{pg_{\mathbf{x}}g_{\mathbf{z}}p_{\mathbf{z}},py_{\mathbf{x}}y_{\mathbf{z}}p_{\mathbf{z}},yg_{\mathbf{x}}y_{\mathbf{z}}g_{\mathbf{z}}\}$ $(m_{1}e_{2}e_{3},m_{2}e_{1}e_{3},m_{3}e_{1}e_{2})$ 1 Elementary III.3.6
$\{(\mathbf{c}_{i})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{k})_{\mathbf{x}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}\}$ $(e_{i},m_{j},m_{k})$ 3 Elementary III.3.1
$\{(\mathbf{c}_{i})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{k})_{\mathbf{x}}(\mathbf{c}_{k})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}(\mathbf{c}_{j})_{\mathbf{z}}\}$ $(e_{i},m_{j}e_{k},m_{k}e_{j})$ 3 Elementary III.3.6
$\{(\mathbf{c}_{i})_{\mathbf{z}},(\mathbf{c}_{j})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}\}$ $(e_{i},e_{j},m_{k})$ 3 Elementary III.3.2
$\{(\mathbf{c}_{i})_{\mathbf{z}}(\mathbf{c}_{j})_{\mathbf{z}},(\mathbf{c}_{j}\mathbf{c}_{k})_{\mathbf{x}}(\mathbf{c}_{i}\mathbf{c}_{k})_{\mathbf{x}},(\mathbf{c}_{k})_{\mathbf{z}}\}$ $(e_{i}e_{j},m_{i}m_{j},e_{k})$ 3 Elementary III.3.3
$\{(\mathbf{c}_{i})_{\mathbf{z}}(\mathbf{c}_{j})_{\mathbf{z}},(\mathbf{c}_{j}\mathbf{c}_{k})_{\mathbf{x}}(\mathbf{c}_{i}\mathbf{c}_{k})_{\mathbf{x}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}\}$ $(e_{i}e_{j},m_{i}m_{j},m_{k})$ 3 Elementary III.3.3
$\{(\mathbf{c}_{i})_{\mathbf{z}}(\mathbf{c}_{j})_{\mathbf{z}},(\mathbf{c}_{j}\mathbf{c}_{k})_{\mathbf{x}}(\mathbf{c}_{i}\mathbf{c}_{k})_{\mathbf{x}}(\mathbf{c}_{k})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}(\mathbf{c}_{j})_{\mathbf{z}}\}$ $(e_{i}e_{j},m_{i}m_{j}e_{k},m_{k}e_{j})$ 3 Elementary III.3.6
$\{y_{\mathbf{z}}g_{\mathbf{z}},g_{\mathbf{z}}p_{\mathbf{z}},pg_{\mathbf{x}}py_{\mathbf{x}}yg_{\mathbf{x}}\}$ $(e_{1}e_{2},e_{2}e_{3},m_{1}m_{2}m_{3})$ 1 Elementary III.3.4
$\{pg_{\mathbf{x}}py_{\mathbf{x}},py_{\mathbf{x}}yg_{\mathbf{x}},y_{\mathbf{z}}g_{\mathbf{z}}p_{\mathbf{z}}\}$ $(m_{1}m_{2},m_{2}m_{3},e_{1}e_{2}e_{3})$ 1 Elementary III.3.5
$\{pg_{\mathbf{x}}py_{\mathbf{x}}y_{\mathbf{z}}g_{\mathbf{z}},py_{\mathbf{x}}yg_{\mathbf{x}}y_{\mathbf{z}},y_{\mathbf{z}}g_{\mathbf{z}}p_{\mathbf{z}}$} $(m_{1}m_{2}e_{1}e_{2},m_{2}m_{3}e_{1},e_{1}e_{2}e_{3})$ 1 Elementary III.3.6
$\{pg_{\mathbf{xs}},py_{\mathbf{xs}},yg_{\mathbf{xs}}\}$ $(m_{1}s_{2,3}^{(2)},m_{2}s_{3,1}^{(2)},m_{3}s_{1,2}^{(2)})$ 1 Magic IV.4

### v4

@@ 4.6.3 The $\{y_{\mathbf{z}},g_{\mathbf{z}},p_{\mathbf{z}}\}$-boundary
4.6.4 The folded boundaries
In this subsection, we investigate the condensation properties of the folded boundaries introduced in Section 3.3.3.
Consider the $\{y_{\mathbf{z}}g_{\mathbf{z}},pg_{\mathbf{x}}py_{\mathbf{x}},yg_{\mathbf{x}}\}$-boundary as an example for the fold|m|m-boundaries. After conjugation of transversal-$T$ gate, the boundary condensation set becomes $\{y_{\mathbf{z}}g_{\mathbf{z}},pg_{\mathbf{xs}}py_{\mathbf{xs}},yg_{\mathbf{x}}\}$. One can check the commutator between $\overline{X_{pg}X_{py}}$ and the boundary $XS$-stabilizer $\mathcal{A}_{y}^{boundary}(XS)\mathcal{A}_{g}^{boundary}(XS)$ is equivalent to the product of $Z$-stabilizers on the boundary. Therefore, we have the following relation:
@@ 4.6.4 The folded boundaries
In the TQFT description, this corresponds to Eq. (55), Eq. (56) and Eq. (58). The total number of distinct types of fold|m|m-boundary is 3.
For the fold|e|e-boundaries, let’s consider the $\{y_{\mathbf{z}}g_{\mathbf{z}},pg_{\mathbf{x}}py_{\mathbf{x}},p_{\mathbf{z}}\}$-boundary as an example. After conjugation of the transversal-$T$ gate, the condensation set becomes $\{y_{\mathbf{z}}g_{\mathbf{z}},pg_{\mathbf{xs}}py_{\mathbf{xs}},p_{\mathbf{z}}\}$. Similarly, one can check the commutator between $\overline{X_{pg}X_{py}}$ and the boundary $XS$-stabilizer $\mathcal{A}_{y}^{boundary}(XS)\mathcal{A}_{g}^{boundary}(XS)$ equals identity within the code space. Therefore, $pg_{\mathbf{xs}}py_{\mathbf{xs}}$ can condense on the bounday, which is equivalent to say
@@
This relation corresponds to Eq. (56). Furthermore, one can check all the three types of $S$-domain walls can condense on this boundary, which corresponds to Eq. (54) in the perspective of TQFT. The total number of distinct types of the fold|e|e-boundary is 3.
4.6.5 The $pg_{\mathbf{x}}py_{\mathbf{x}}yg_{\mathbf{x}}$- and the $y_{\mathbf{z}}g_{\mathbf{z}}p_{\mathbf{z}}$-boundaries
Consider the lattice model we introduce in Section 3.3.4 and 3.3.5. After the conjugation of transversal-$T$ gate, the boundary condensation set becomes $\{y_{\mathbf{z}}g_{\mathbf{z}},g_{\mathbf{z}}p_{\mathbf{z}},pg_{\mathbf{xs}}py_{\mathbf{xs}}yg_{\mathbf{xs}}\}$ and $\{pg_{\mathbf{xs}}py_{\mathbf{xs}},py_{\mathbf{xs}}yg_{\mathbf{xs}},y_{\mathbf{z}}g_{\mathbf{z}}p_{\mathbf{z}}\}$, respectively.
@@
$\{y_{\mathbf{z}}g_{\mathbf{z}},g_{\mathbf{z}}p_{\mathbf{z}},pg_{\mathbf{xs}}py_{\mathbf{xs}}yg_{\mathbf{xs}}\}$
@@
4.7 The nested boundaries
Figure 9: An illustration of the $S$ $(\mathrm{CZ}$)-domain wall sweeping across a codimension-1 submanifold in the presence of an boundary with at least one $Z$-type condensation. (a) When the boundary has one $Z$-type condensations, for example $y_{\mathbf{z}}$, the $S$-domain wall on the $pg$-membrane cannot condense on it. (b) However, the other two $S$-domain walls on the $py$- and ygyg-membranes can condense on this boundary. (c) When the boundary has two or more $Z$-type condensations, all of the three types of $S$-domain walls can condense on it.
As suggested in Section 4.2, there are four types of boundaries that some of the $S$-domain walls cannot condense on, i.e., $(e_{i},m_{j},m_{k})$, $(m_{1},m_{2},m_{3})$, $(e_{i}e_{j},m_{i}m_{j},m_{k})$ and $(m_{1}m_{2},m_{2}m_{3},e_{1}e_{2}e_{3})$ boundaries in the TQFT descriptions. The corresponding color-code boundaries are the $\mathbf{c}_{i}$-color boundaries, the $X$-boundary, the fold|m|m-boundaries, and the $y_{\mathbf{z}}g_{\mathbf{z}}p_{\mathbf{z}}$-boundary, respectively. We study the properties of these boundaries in this subsection, and summarize them in Fig. 10.
4.7.1 The $\mathbf{c}_{i}$-color boundary
Consider the $\{(\mathbf{c}_{i})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}},(\mathbf{c}_{i}\mathbf{c}_{k})_{\mathbf{x}}\}$-boundary, the $S_{\mathbf{c}_{i}\mathbf{c}_{j}}$- or $S_{\mathbf{c}_{i}\mathbf{c}_{k}}$-domain wall can condense on the boundary. However, the $S_{\mathbf{c}_{j}\mathbf{c}_{k}}$-domain wall can not condense on it. It attaches to the boundary and generates a codimension-2 (1D) domain-wall defect. The condensation set on that defect becomes $\{(\mathbf{c}_{i})_{\mathbf{z}},(\mathbf{c}_{k})_{\mathbf{z}}(\mathbf{c}_{i}\mathbf{c}_{k})_{\mathbf{x}},(\mathbf{c}_{j})_{\mathbf{z}}(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}\}$, while other part of the boundary is unchanged. The corresponding condensation in the TQFT description is $(e_{i},m_{j}e_{k},m_{k}e_{j})$, which has been introduced in Section 4.2. The total number of distinct types of the $\mathbf{c}_{i}$-color boundaries with non-condensing $S$-domain wall is 3. We denote this type of boundary as the nested boundary, as illustrated in Fig. 8, Fig. 9(a), and Fig. 10(d).
4.7.2 The $X$-boundary
@@
4.7.3 The fold|m|m-boundary
Consider the $\{y_{\mathbf{z}}g_{\mathbf{z}},pg_{\mathbf{x}}py_{\mathbf{x}},yg_{\mathbf{x}}\}$-boundary as an example. the $S_{yg}$ domain wall can condense on it. However, the $S_{pg}$ and $S_{yp}$ domain wall cannot. At the intersection of this boundary and the $S_{pg}$-domain wall, the condensation set becomes $\{y_{\mathbf{z}}g_{\mathbf{z}},p_{\mathbf{z}}pg_{\mathbf{x}}py_{\mathbf{x}},g_{\mathbf{z}}yg_{\mathbf{x}}\}$, while the other part of the boundary remain unchanged. In the TQFT description, this boundary is denoted as $(e_{1}e_{2},e_{3}m_{1}m_{2},e_{2}m_{3})$. The total number of distinct types of fold|m|m-boundaries with nested boundaries is 3, as illustrated in Fig. 10(c).
@@
4.7.4 The $y_{\mathbf{z}}g_{\mathbf{z}}p_{\mathbf{z}}$-boundary
The condensation set on the $y_{\mathbf{z}}g_{\mathbf{z}}p_{\mathbf{z}}$-boundary is $\{pg_{\mathbf{x}}py_{\mathbf{x}},py_{\mathbf{x}}yg_{\mathbf{x}},y_{\mathbf{z}}g_{\mathbf{z}}p_{\mathbf{z}}\}$. The lattice construction is given in Section 3.3.5. One can show that by sweeping the $S$-domain wall across the codimension-1 submanifold, at the intersection with the boundary, the $S$-domain wall cannot terminate. In the TQFT description, this relation corresponds to Eq. (59). The total number of distinct types of nested boundaries associated with this boundary is 1, as illustrated in Fig. 10(b).
4.7.5 The other boundaries
@@
4.8 Summary of boundary types
In summary, we identify 101 distinct types of boundaries for the 3D color code, with one being the magic boundary and 70 being nested boundaries. While a trivial color permutation ($y\to g$, $g\to p$, $p\to r$, and $r\to y$) can generate additional boundary types, the new boundaries created by this permutation exhibit identical physical properties, differing only by relabeling of the lattice and excitations. Therefore, we exclude these from our count. We summarize the properties of these boundaries in Table 1, Table 2, and Fig. 10.
@@ 4.8 Summary of boundary types
$\{y_{\mathbf{z}},g_{\mathbf{z}},p_{\mathbf{z}}\}$ $(e_{1},e_{2},e_{3})$ 1 Elementary 3.1
$\{pg_{\mathbf{x}},py_{\mathbf{x}},yg_{\mathbf{x}}\}$ $(m_{1},m_{2},m_{3})$ 1 Elementary 3.2
$\{(\mathbf{c}_{j}\mathbf{c}_{k})_{\mathbf{x}}(\mathbf{c}_{j})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{k})_{\mathbf{x}}(\mathbf{c}_{i})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}\}$ $(m_{i}e_{j},m_{j}e_{i},m_{k})$ 3 Elementary 3.3.6
$\{(\mathbf{c}_{j}\mathbf{c}_{k})_{\mathbf{x}}(\mathbf{c}_{j})_{\mathbf{z}}(\mathbf{c}_{k})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{k})_{\mathbf{x}}(\mathbf{c}_{i})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}(\mathbf{c}_{i})_{\mathbf{z}}\}$ $(m_{i}e_{j}e_{k},m_{j}e_{i},m_{k}e_{i})$ 3 Elementary 3.3.6
$\{pg_{\mathbf{x}}g_{\mathbf{z}}p_{\mathbf{z}},py_{\mathbf{x}}y_{\mathbf{z}}p_{\mathbf{z}},yg_{\mathbf{x}}y_{\mathbf{z}}g_{\mathbf{z}}\}$ $(m_{1}e_{2}e_{3},m_{2}e_{1}e_{3},m_{3}e_{1}e_{2})$ 1 Elementary 3.3.6
$\{(\mathbf{c}_{i})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{k})_{\mathbf{x}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}\}$ $(e_{i},m_{j},m_{k})$ 3 Elementary 3.3.1
$\{(\mathbf{c}_{i})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{k})_{\mathbf{x}}(\mathbf{c}_{k})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}(\mathbf{c}_{j})_{\mathbf{z}}\}$ $(e_{i},m_{j}e_{k},m_{k}e_{j})$ 3 Elementary 3.3.6
$\{(\mathbf{c}_{i})_{\mathbf{z}},(\mathbf{c}_{j})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}\}$ $(e_{i},e_{j},m_{k})$ 3 Elementary 3.3.2
$\{(\mathbf{c}_{i})_{\mathbf{z}}(\mathbf{c}_{j})_{\mathbf{z}},(\mathbf{c}_{j}\mathbf{c}_{k})_{\mathbf{x}}(\mathbf{c}_{i}\mathbf{c}_{k})_{\mathbf{x}},(\mathbf{c}_{k})_{\mathbf{z}}\}$ $(e_{i}e_{j},m_{i}m_{j},e_{k})$ 3 Elementary 3.3.3
$\{(\mathbf{c}_{i})_{\mathbf{z}}(\mathbf{c}_{j})_{\mathbf{z}},(\mathbf{c}_{j}\mathbf{c}_{k})_{\mathbf{x}}(\mathbf{c}_{i}\mathbf{c}_{k})_{\mathbf{x}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}\}$ $(e_{i}e_{j},m_{i}m_{j},m_{k})$ 3 Elementary 3.3.3
$\{(\mathbf{c}_{i})_{\mathbf{z}}(\mathbf{c}_{j})_{\mathbf{z}},(\mathbf{c}_{j}\mathbf{c}_{k})_{\mathbf{x}}(\mathbf{c}_{i}\mathbf{c}_{k})_{\mathbf{x}}(\mathbf{c}_{k})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}(\mathbf{c}_{j})_{\mathbf{z}}\}$ $(e_{i}e_{j},m_{i}m_{j}e_{k},m_{k}e_{j})$ 3 Elementary 3.3.6
$\{y_{\mathbf{z}}g_{\mathbf{z}},g_{\mathbf{z}}p_{\mathbf{z}},pg_{\mathbf{x}}py_{\mathbf{x}}yg_{\mathbf{x}}\}$ $(e_{1}e_{2},e_{2}e_{3},m_{1}m_{2}m_{3})$ 1 Elementary 3.3.4
$\{pg_{\mathbf{x}}py_{\mathbf{x}},py_{\mathbf{x}}yg_{\mathbf{x}},y_{\mathbf{z}}g_{\mathbf{z}}p_{\mathbf{z}}\}$ $(m_{1}m_{2},m_{2}m_{3},e_{1}e_{2}e_{3})$ 1 Elementary 3.3.5
$\{pg_{\mathbf{x}}py_{\mathbf{x}}y_{\mathbf{z}}g_{\mathbf{z}},py_{\mathbf{x}}yg_{\mathbf{x}}y_{\mathbf{z}},y_{\mathbf{z}}g_{\mathbf{z}}p_{\mathbf{z}}$} $(m_{1}m_{2}e_{1}e_{2},m_{2}m_{3}e_{1},e_{1}e_{2}e_{3})$ 1 Elementary 3.3.6
$\{pg_{\mathbf{xs}},py_{\mathbf{xs}},yg_{\mathbf{xs}}\}$ $(m_{1}s_{2,3}^{(2)},m_{2}s_{3,1}^{(2)},m_{3}s_{1,2}^{(2)})$ 1 Magic 4.5